# Differentially Private Distributed Inference

### v6

@@ 4.3 Differentially Private Distributed Online Learning from Intermittent Streams
iterations, where $l_{n,\Theta}$ and $Q_{n,\Theta}$ are properties of the signal structure defined in Equations 9 and 10, running Algorithm 3 yields $\mathbb{P}\left[\theta^{\circ}=\hat{\theta}^{\mathrm{OL}}_{i,T}\right]\geq 1-\eta$ as long as $\sum_{i=1}^{n}\xi_{i}D_{KL}\left(\ell_{i}(\cdot|\bar{\theta})|\ell_{i}(\cdot|\theta^{\circ})\right)>0$ for all $\bar{\theta}\neq\theta^{\circ}$. Subsequently, the noise distributions that optimize runtime are $\mathcal{D}_{i,t}^{\star}=\mathrm{Lap}\left({\Delta_{n,\Theta}|\Theta|}/{{\varepsilon}}\right)$. The resulting estimates are ${\varepsilon}$-DP with respect to the private signals.
Also, similarly to distributed MLE algorithms, the algorithm adds noise to each of the $|\Theta|$ states, and an attacker would have more observations of the same signals; therefore, we need to rescale the privacy guarantee of each run and each state by $|\Theta|$.
@@ 5 Simulation Study
Runtime study. In Section 14.2, we characterize the computational complexity of our algorithms, running in polynomial time and space. In Section 14.3, we perform a runtime study on our proposed algorithms and show that our algorithm can run between $\sim 10^{-2}$ and $10$ seconds for values of $n$ ranging from $n=3$ to $n=96$, which is significantly faster (up to 1000×1000$\times$1000 ×) than existing methods that rely on homomorphic encryption (Froelicher et al. 2021). Fully homomorphic encryption and multiparty computation are cryptographic techniques that allow computation on encrypted data (by single or multiple severs, respectively) without the need to access decrypted data. The strong cryptographic guarantees of these methods come at a high cost to computational efficiency, and their implementation requires complex coordination and information technology (IT) infrastructure across multiple institutions.
@@ 6 Discussion
    The AIDS Clinical Trials Dataset was obtained from Kaggle: https://www.kaggle.com/datasets/tanshihjen/aids-clinical-trials (Accessed on February 14, 2025). Original information about the clinical trial can be found at https://clinicaltrials.gov/study/NCT00000625 (accessed February 14, 2025).
@@
The codes for reproducing the simulation results and figures are open-source and can be accessed at https://github.com/papachristoumarios/dp-social-learning.
@@ 7 Related Work
Privacy-preserving methods in healthcare contexts. Data privacy and security are critical in healthcare contexts (ARPA-H 2024). An example is multicenter clinical trials, which are often difficult to coordinate and conduct; see, for example, the AIDS Clinical Trials Group (Campbell et al. 2012, Hammer et al. 1996). In the literature, a variety of methods and protocols have been proposed for privacy-preserving data sharing between collaborating healthcare centers and data providers; see, for example, Wan et al. (2022), Rehm et al. (2021), Froelicher et al. (2021), Munjal and Bhatia (2023). Most of these methods are based on holomorphic encryption schemes and rely on centralized processing authority; see, e.g., Froelicher et al. (2021), Munjal and Bhatia (2023), Geva et al. (2023)). In contrast, our method allows for fully distributed computation and does not require a centralized party to coordinate communication or computations between the centers. Furthermore, our method is based on statistical disclosure control following the differential privacy framework, which offers significant advantages in runtime and flexibility to accommodate a variety of statistical models, including survival analysis (Geva et al. 2023, Froelicher et al. 2021), logistic regression (Geva et al. 2023), and linear regression (Froelicher et al. 2021).
@@ 9.1 Log-Belief Ratio Notations
where $\Gamma_{n,\Theta}=\max_{i\in[n],\hat{\theta}\in\Theta}|\log\bm{\gamma}_{i}(\hat{\theta})|$ and $|\lambda_{2}(A+I)|/2<1$ is the SLEM of $(A+I)/2$; $A$ is doubly stochastic so $20<|\lambda_{n}(A+I)|\leq\dots\leq|\lambda_{n-1}(A+I)|<\lambda_{1}(A+I)=$ 2. Note that the right-hand side in Equation 19 goes to 00 as $t\to\infty$, which implies that (by Markov’s inequality) $\overline{\bm{\psi}}_{k,t}(\hat{\theta},\check{\theta})\overset{L_{2}}{\to}\frac{2^{t}}{n}Z_{k}(\hat{\theta},\check{\theta})\mathds{1}$.
@@ 12 Sensitivity of the Proportional Hazards Model
We derive the sensitivity of the proportional hazards model used in our experiments. Specifically, each center has $n_{i}$ data points $\bm{S}_{i}$, and the treatment variable is bounded above by $1$ (that is, $\max_{1\leq j\leq n_{i}}|\bm{x}_{i,j}|\leq B_{x}$). The state $\theta$ obeys $|\theta|\leq B_{\theta}$. Let $\bm{S}_{i}^{\prime}$ be a data set of $n_{i}$ points such that $\bm{S}_{i}$ and $\bm{S}_{i}^{\prime}$ differ in patient $k$. The $\log$ partial likelihood can be written as
$\log\ell_{i}(\bm{S}_{i}|\theta)=\sum_{j=1}^{n_{i}}\bm{\delta}_{ij}\left(\theta\bm{x}_{ij}-\log\sum_{r\in R(j)}e^{\theta\bm{x}_{ik}}\right)=\sum_{j=1}^{n_{i}}\bm{\delta}_{ij}\theta\bm{x}_{ij}-\sum_{j=1}^{n_{i}}\bm{\delta}_{ij}\log\sum_{r\in R(j)}e^{\theta\bm{x}_{ir}}$
@@
  • If $k$ is removed and $k$ in some risk sets, removing $k$ affects the denominator $\sum_{r\in R(j)}e^{\theta\bm{x}_{ir}}$ for all $R(j)$ such that $k\in R(j)$.
@@
Thus, the term $\theta\bm{x}_{ij}$ in the numerator is always bounded by $B_{\theta}B_{x}$ due to Hölder’s inequality. The denominator $\sum_{r\in R(j)}e^{\theta\bm{x}_{ir}}$ is a sum of exponentials, and its change is bounded by $e^{B_{\theta}B_{x}}$ and its logarithm by $B_{\theta}B_{x}$. Therefore, $\Delta_{n,\Theta}=2B_{\theta}B_{x}$.
@@ 13 Lower Bounds for Distributed Hypothesis Testing
In this section, we devise information-theoretic lower bounds that are applicable to any belief-exchange algorithms. We focus on the hypothesis testing scenario with $\Theta$ consisting of a null hypothesis ($\theta=0$) and an alternative hypothesis ($\theta=1$). Initially, each agent has access to a model $\ell_{i}(\cdot|\theta)$ and an ${\varepsilon}$-DP privacy mechanism $\mathcal{M}_{i}$ which the agent uses $K$ times to construct a private signal $\bm{y}_{i,k}=\mathcal{M}_{i}(\bm{s}_{i,k})$. Agents can exchange the private signals $\bm{y}_{i,k}$ that they possess in each iteration $T$, and have access to a statistical test $\mathcal{A}:\mathcal{R}^{nK}\to\{0,1\}$ that, given private signals $\bm{y}_{1,1},\dots,\bm{y}_{n,K}$, outputs 1 if $\theta=0$ is rejected and 0 if we fail to reject $\theta=0$. We want to find a lower bound on communication complexity, $K\cdot T$, assuming that $\mathcal{A}$ achieves a Type I error rate of $\alpha$ and a Type II error rate of $1-\beta$. To devise the lower bound, we rely on classic results from the bandits literature (Slivkins et al. 2019).
@@ 15 Table of Notations
$\theta^{\star}$, $\bar{\theta}$ MLE, non-MLE state
@@
$l_{n,\Theta}$ minimum KL-divergence between any non-MLE state and an MLE state

### v7

@@ 4.3 Differentially Private Distributed Online Learning from Intermittent Streams
iterations, where $l_{n,\Theta}$ and $Q_{n,\Theta}$ are properties of the signal structure defined in Equations 10 and 11, running Algorithm 3 yields $\mathbb{P}\left[\theta^{\circ}=\hat{\theta}^{\mathrm{OL}}_{i,T}\right]\geq 1-\eta$ as long as $\sum_{i=1}^{n}\xi_{i}D_{KL}\left(\ell_{i}(\cdot|\bar{\theta})|\ell_{i}(\cdot|\theta^{\circ})\right)>0$ for all $\bar{\theta}\neq\theta^{\circ}$. Subsequently, the noise distributions that optimize runtime are $\mathcal{D}_{i,t}^{\star}=\mathrm{Lap}\left({\Delta_{n,\Theta}|\Theta|}/{{\varepsilon}}\right)$. The resulting estimates are ${\varepsilon}$-DP with respect to the private signals.
Note that similarly to distributed MLE algorithms, here we also need to scale the privacy budget by $|\Theta|$ to account for the fact that the algorithm accesses the private signals and adds noise for each of the $|\Theta|$ states, providing an attacker with as many observations of the private signals.
@@ 5 Simulation Study
Runtime study. In Section 14.2, we characterize the computational complexity of our algorithms, running in polynomial time and space. In Section 14.3, we perform a runtime study on our proposed algorithms and show that our algorithm can run between $\sim 10^{-2}$ and $10$ seconds for values of $n$ ranging from $n=3$ to $n=96$, which is significantly faster (up to 1000×1000$\times$1000 ×) than existing methods that rely on homomorphic encryption (Froelicher et al. 2021). Fully homomorphic encryption and multiparty computation are cryptographic techniques that allow computation on encrypted data (by single or multiple severs, respectively) without the need to access decrypted data. The strong cryptographic guarantees of these methods come at a high cost to computational efficiency, and their implementation requires information technology (IT) infrastructure and large-scale, high-performance computational resources across distributed sites, which are beyond the technical capabilities of many organizations.
@@ 6 Discussion
    The AIDS Clinical Trials data set was obtained from Kaggle: https://www.kaggle.com/datasets/tanshihjen/aids-clinical-trials (accessed February 14, 2025). Original information about the clinical trial can be found at https://clinicaltrials.gov/study/NCT00000625 (accessed February 14, 2025).
@@
The codes for reproducing the simulation results and figures are open source and can be accessed at https://github.com/papachristoumarios/dp-social-learning.
@@ 7 Additional Related Work
Privacy-preserving methods in healthcare. Privacy and data security are critical in healthcare contexts (ARPA-H 2024). An example is multicenter clinical trials, which are often difficult to coordinate and conduct; see, for example, the AIDS Clinical Trials Group (Campbell et al. 2012, Hammer et al. 1996). In the literature, a variety of methods and protocols have been proposed for privacy-preserving data sharing between collaborating healthcare centers and data providers; see, for example, Wan et al. (2022), Rehm et al. (2021), Froelicher et al. (2021), Munjal and Bhatia (2023). Most of these methods are based on holomorphic encryption schemes and rely on centralized processing authority; see, e.g., Froelicher et al. (2021), Munjal and Bhatia (2023), Geva et al. (2023)). In contrast, our method allows for fully distributed computation and does not require a centralized party to coordinate communication or computations between the centers. Furthermore, our method is based on statistical disclosure control following the differential privacy framework, which offers significant advantages in runtime and flexibility to accommodate a variety of statistical models, including survival analysis (Geva et al. 2023, Froelicher et al. 2021), logistic regression (Geva et al. 2023), and linear regression (Froelicher et al. 2021).
@@ 9.1 Log-Belief Ratio Notations
where $\Gamma_{n,\Theta}=\max_{i\in[n],\hat{\theta}\in\Theta}|\log\bm{\gamma}_{i}(\hat{\theta})|$ and $|\lambda_{2}(A+I)|/2<1$ is the SLEM of $(A+I)/2$, henceforth also denoted by $a^{\star}_{n}$; $A$ is doubly stochastic so $20<|\lambda_{n}(A+I)|\leq|\lambda_{n-1}(A+I)|\leq\dots\leq|\lambda_{2}(A+I)|<\lambda_{1}(A+I)=$ 2. Note that the right-hand side in Equation 20 goes to 00 as $t\to\infty$, which implies that (by Markov’s inequality) $\overline{\bm{\psi}}_{k,t}(\hat{\theta},\check{\theta})\overset{L_{2}}{\to}\frac{2^{t}}{n}Z_{k}(\hat{\theta},\check{\theta})\mathds{1}$.
@@ 12 Sensitivity of the Proportional Hazards Model
We derive the sensitivity of the proportional hazards model used in our experiments. Specifically, each center has $n_{i}$ data points $\bm{S}_{i}$, and the treatment variable is bounded above by $1$ (that is, $\max_{1\leq j\leq n_{i}}|\bm{x}_{i,j}|\leq B_{x}=1$). The state $\theta$ obeys $|\theta|\leq B_{\theta}$. Let $\bm{S}_{i}^{\prime}$ be a data set of $n_{i}$ points such that $\bm{S}_{i}$ and $\bm{S}_{i}^{\prime}$ differ in patient $k$. The $\log$ partial likelihood can be written as
$\log\ell_{i}(\bm{S}_{i}|\theta)=\log\prod_{j:\delta_{ij}=1}\frac{e^{\theta\bm{x}_{ij}}}{\sum_{r\in R(j)}e^{\theta\bm{x}_{ir}}}=\sum_{j=1}^{n_{i}}\bm{\delta}_{ij}\left(\theta\bm{x}_{ij}-\log\sum_{r\in R(j)}e^{\theta\bm{x}_{ir}}\right)=\sum_{j=1}^{n_{i}}\bm{\delta}_{ij}\theta\bm{x}_{ij}-\sum_{j=1}^{n_{i}}\bm{\delta}_{ij}\log\sum_{r\in R(j)}e^{\theta\bm{x}_{ir}}$
@@
  • If $k$ is removed and $k$ is in some risk sets, removing $k$ affects the denominator $\sum_{r\in R(j)}e^{\theta\bm{x}_{ir}}$ for all $R(j)$ such that $k\in R(j)$.
@@
Thus, the term $e^{\theta\bm{x}_{ij}}$ in the numerator is always bounded by $e^{B_{\theta}B_{x}}$ and its logarithm by $B_{\theta}B_{x}$. The denominator $\sum_{r\in R(j)}e^{\theta\bm{x}_{ir}}$ is a sum of exponentials, and its change is bounded by $e^{B_{\theta}B_{x}}$ and its logarithm by $B_{\theta}B_{x}$. Therefore, $\Delta_{n,\Theta}\leq 2B_{\theta}B_{x}=2B_{\theta}$.
@@ 13 Lower Bounds for Distributed Hypothesis Testing
In this section, we devise information-theoretic lower bounds that are applicable to any belief-exchange algorithms. We focus on the hypothesis testing scenario with $\Theta$ consisting of a null hypothesis ($\theta=0$) and an alternative hypothesis ($\theta=1$). Initially, each agent has access to a model $\ell_{i}(\cdot|\theta)$ and an ${\varepsilon}$-DP privacy mechanism $\mathcal{M}_{i}$ which the agent uses $K$ times to build a private signal $\bm{y}_{i,k}=\mathcal{M}_{i}(\bm{s}_{i,k})$. Agents can exchange the private signals $\bm{y}_{i,k}$ that they possess in each iteration $T$, and have access to a statistical test $\mathcal{A}:\mathcal{R}^{nK}\to\{0,1\}$ that, given private signals $\bm{y}_{1,1},\dots,\bm{y}_{n,K}$, outputs 1 if $\theta=0$ is rejected and 0 if we fail to reject $\theta=0$. We want to find a lower bound on communication complexity, $K\cdot T$, assuming that $\mathcal{A}$ achieves a Type I error rate of $\alpha$ and a Type II error rate of $1-\beta$. To devise the lower bound, we rely on classic results from the bandits literature (Slivkins et al. 2019).
@@ 15 Table of Notations
$\theta^{\star}$, $\bar{\theta}$, $\theta^{\circ}$ MLE, non-MLE/false state, true state (in the online learning case)
@@
$l_{n,\Theta}$ minimum KL-divergence between non-MLE and MLE states (false and true states in the OL case)